On a research it was found that 47% of residents travelers were women and 53% men. A random sample of 500 travelers on a large airline revealed that of those 500, 263 were women. Using a=0.05 compute CV, and TV, and state the hypotheses and your conclusion.


1
Expert's answer
2021-05-07T10:35:27-0400

Probability of women residents "p=0.47"

Also, 263 were women from sample of 500 , "\hat{p}=\dfrac{263}{500}=0.526"


Let "H_o:p=\hat{p}" be the Null hypothesis.

"H_a:p\neq \hat{p}" be the alternate hypothesis.


"\alpha=0.05,"

"z_{\frac{\alpha}{2}}=z_{0.025}=1.96"


Critical value is 1.96.


Calculating Test value(TV)-

"z=\dfrac{\hat{p}-p}{\sqrt{\frac{p(1-p)}{n}}}"


"=\dfrac{0.526-0.47}{\sqrt{\frac{0.47(1-0.47)}{500}}}=\dfrac{0.056}{0.02232}=2.5089"

Conclusion: As the test value is greater than the critical value, So Null Hypotesis is rejected i.e "p\neq \hat{p}"